OpenStudy (anonymous):

integrate 2/square roots x -1

OpenStudy (anonymous):

\[2/\sqrt{x-1}\]

OpenStudy (anonymous):

Use the following information to rewrite: \[\sqrt{x}=x^{1/2}\] \[1/x^2=x^{-2}\]

OpenStudy (kropot72):

By inspection the result is: \[4/3(x-1)^{3/2}+C\] This can be checked by differentiating back to the original expression to be integrated>

OpenStudy (anonymous):

i am sorry. it's (2/square roots x) -1

OpenStudy (anonymous):

is the answer (4/square roots x) -1?

OpenStudy (kropot72):

Is this it? \[\int\limits_{}^{}(2/\sqrt{x})-1.dx\]

OpenStudy (anonymous):

Umm whenever you integrate a constant the result will be that constant times the variable you integrated from. In @kropot72 integral you can split it into two parts the second part will consist of the integral of -1 which results in -1x

OpenStudy (kropot72):

The answer is as follows:\[4\sqrt{x}-x+C\]

OpenStudy (anonymous):

The first part you have to integrate \[2/\sqrt{x}\] in respect to x.

OpenStudy (anonymous):

THANKYOU